# Conformation and dynamics of wet tangentially-driven active filaments

### v1

@@ I Introduction
Active particles, driven by self-generated mechanisms, have emerged as a fascinating research area as they exhibit novel patterns of non-equilibrium self-organization with no counterparts in equilibrium systems. Active systems, encompassing a diverse range of entities such as bird flocks, fish schools, swimming bacteria, and artificial microswimmers, offer valuable insights into the governing principles of non-equilibrium self-organization in active matter Vicsek and Zafeiris (2012); Bechinger et al. (2016). Recent research interests have extended beyond point-like or rigid active particles (e.g., active Brownian particles) to include particles with complex shapes or internal degrees of freedom. Prominent examples include active polymers and filaments Winkler et al. (2017); Winkler and Gompper (2020), which have recently attracted a lot of attention.
@@
Active filaments and polymers can be found on a wide range of scales from ATP-powered chromatin in cell nucleus and motor-driven actin and microtubules to microscopic worms and snakes. Inspired by experiments of active biological filaments a diverse range of theoretical and computational models for active polymers have been proposed in recent years Winkler et al. (2017); Winkler and Gompper (2020) including scalar Smrek and Kremer (2018); Chubak et al. (2020) and force-driven active polymer models where the active force can be stochastic Eisenstecken et al. (2016); Ghosh and Gov (2014), tangential Isele-Holder et al. (2015); Bianco et al. (2018), perpendicular Prathyusha et al. (2022) to the backbone. Among these, the tangentially-driven active polymer model has garnered increased interest due to its relevance to filament-like systems, such as motor-driven actin and microtubule filaments, elongated bacteria like Proteus mirabilis Lough et al. (2023), and worm species such as T. tubifex  Deblais et al. (2020) that propel themselves along their backbone.
@@
Microscopic active filament-like systems, such as motor-driven biopolymers and bacteria swim in aquatic environments. Consequently, their shape and dynamics are influenced not only by their intrinsic activity mechanisms but also by fluid-mediated interactions. It is well known that the dynamics of passive polymer solutions are significantly affected by hydrodynamic interactions Zimm (1956); Doi et al. (1988); Nikoubashman et al. (2016). Therefore, it is important to elucidate the interplay between internal degrees of freedom, activity, and hydrodynamic interactions on the conformation and dynamics of active polymers. Integrating shape, activity, and hydrodynamic interactions in studies of active polymer models not only provides a deeper understanding of biophysical processes involving wet active polymers but also inspires the design of artificial filamentous microswimmers with tailored properties.
@@
The majority of studies on active polymers have ignored hydrodynamic interactions and have focused on the interplay between internal dynamics and activity mechanisms. There are a few studies that investigate the role of hydrodynamic interactions on the structural and dynamical properties of active filaments Jayaraman et al. (2012); Laskar and Adhikari (2015); De Canio et al. (2017); Saintillan et al. (2018); Martín-Gómez et al. (2019, 2019); Mahajan and Saintillan (2022); Stein et al. (2021); Lough et al. (2023). These include studies of active Brownian polymers in solution Martín-Gómez et al. (2019, 2019) and active polymer models where monomers are subject to permanent or stochastic active force dipoles along their backbone Jayaraman et al. (2012); Mahajan and Saintillan (2022). These investigations reveal the significance of hydrodynamic interactions on polymer dynamics and even on conformational properties.
@@
In this work, we analyze the conformational and dynamical properties of tangentially-driven active polymers Isele-Holder et al. (2015) embedded in a fluid where the monomers are externally actuated, focusing on unconfined dilute active polymer solutions. The dry version of this model has been subject of intensive studies Isele-Holder et al. (2015); Fazelzadeh et al. (2023); Bianco et al. (2018); Mokhtari and Zippelius (2019); Karan et al. (2024); Khalilian et al. (2024); Duman et al. (2018); Prathyusha et al. (2018); Vatin et al. (2024). This model is relevant for biological filaments actuated by molecular motors and chains of Janus active colloids bonded by electrostatic or magnetic dipolar interactions Yan et al. (2016); Nariaki and C. Patrick (2020); Chao et al. (2024) resulting in an active force tangential to the backbone. Comparing the conformation and dynamics of dry and wet tangentially-driven polymers we address two important questions. (i) Do hydrodynamic interactions lead to the emergence of new modes of conformation and transport of active polymers? (ii) How do the longtime dynamics of polymers wet active polymers differ from that of dry active polymers?
@@
To account for hydrodynamic interactions, we employ the explicit fluid simulation method of multi particle collision dynamics (MPC). MPC is a mesoscale particle-based simulation method for simulation of fluid, which which naturally incorporates thermal fluctuations. MPC method is able to capture the hydrodynamic correlations correctly Malevanets and Kapral (1999); Howard et al. (2019); Tao et al. (2008); Llahí et al. (2022); Huang et al. (2012) and is widely used for simulating dynamics of colloidal and polymeric solutions. MPC method allows to faithfully reproduce the fluid-mediated interactions with significantly less computational costs. In the context of active systems, it has been utilized to study the dynamics of microswimmers such as generic squirmer model Babu and Stark (2012); Hu et al. (2015); Downton and Stark (2009); Götze and Gompper (2010) and wet active Brownian polymer models Martín-Gómez et al. (2019, 2019). In this article, we present an implementation of wet tangentially-driven polymers in multiparticle collision dynamics (MPC). The polymer model consists of linear bead-spring chains where each monomer experiences an external active force tangential to the local tangent of the backbone Isele-Holder et al. (2015).
@@
The remainder of this article is organized as follows. In section II, we first introduce our simulation setup including the implementation of tangentially actuated polymer model in the MPC fluid, providing definition and values of the relevant dimensionless parameters. In Sec. III, We investigate the effects of hydrodynamic interactions (HI) on conformational properties of active polymers. In Sec. IV, we discuss the shrinkage mechanism induced by fluid-mediated interactions. Then, we present the results on impact of hydrodynamic interactions on orientational and translational dynamics of active polymers. We end our paper with a summary of our most important findings and concluding remarks in Sec. V.
@@ II.1 Multiparticle Collision Dynamics simulations
where $\mathbf{\mathcal{R}}(\alpha)$ is a rotation matrix and $\mathbf{\bm{v}}_{cm}=\frac{1}{N_{c}}\sum^{N_{c}}_{i=1}\mathbf{\bm{v}}_{i}$ is the centre of mass velocity of a collision cell. The updated velocities $\mathbf{\bm{v}}^{\prime}_{i}(t+\delta t)$ are then used for the repeating the streaming step described by Eq. 1 .
@@ II.2 Active polymer model
We model a semiflexible polymers by a bead-spring chain composed of $N_{m}$ active monomers of mass $M$, with positions $\mathbf{\bm{R}}_{i}$, and velocities $\mathbf{\bm{V}}_{i}$ $(i=1,...,N_{m})$. The monomers are connected by harmonic springs described by potential $U_{s}$. The degree of flexibility of polymers is incorporated by restrictions of bond orientations via a bending potential $U_{b}$. The self-avoidance is ensured via excluded-volume interactions $U_{EV}$. The dynamics of each monomer is described by the following equation of motion
@@
$U_{EV}(R_{ij})=\begin{cases}4\epsilon\Bigg{[}\bigg{(}\frac{\sigma}{R_{ij}}\bigg{)}^{12}\bigg{(}\frac{\sigma}{R_{ij}}\bigg{)}^{6}\Bigg{]}+\epsilon\text{,}\quad 0\leq R_{ij}\leq 2^{1/6}\sigma\\ 0\text{,}\quad R_{ij}>2^{1/6}\sigma,\end{cases}$ (7)
$\sigma$ effectively sets the diameter of the monomers, and $R_{ij}=|\mathbf{\bm{R}}_{i}-\mathbf{\bm{R}}_{j}|$ is the distance between the $i$th and $j$th monomers. The active force $\mathbf{\bm{f}}^{a}_{i}$ in Eq. (3) is an external tangential force directed along the backbone of the polymer mimicking the effect of the tangential force exerted by the molecular motors on the polymer backbone. The active force on each bead, except for the end monomers, is modeled as Isele-Holder et al. (2015):
@@ II.3 Coupling of polymer dynamics to fluid
where $\mathbf{\bm{F}}^{a}_{tot}$ is the sum of all the tangential forces applied to the monomers. The sum of forces induced by intramolecular interactions in the polymer cancel out due to Newton’s third law, thus we can focus purely on the total active force $\mathbf{\bm{F}}^{a}_{tot}$ itself. In confined systems, the global fluid flow creates a backflow. We can achieve a similar backflow in systems with periodic boundary conditions by applying an additional backflow force to all monomers and MPC particles so that the total force becomes null, i.e. $\mathbf{\bm{F}}^{a}_{tot}(t)+\mathbf{\bm{F}}^{b}_{tot}(t)$. We distribute the backflow forces acting on the monomers $\mathbf{\bm{F}}^{b}_{mon}$ and MPC particles $\mathbf{\bm{F}}^{b}_{MPC}$ according to their masses Martin-Gomez et al. (2020); Llahí et al. (2022) as
@@
Here, $M$ and $m$ are the masses of the monomers and MPC particles and $N_{m}$ and $N$ denote the number of monomers and MPC particles respectively. The magnitude of the backflow force is governed by the ratio between the particle mass and the total mass of the particles in the system. Assuming that the timescale of changes of polymer conformation are larger than the collision time $\delta t$, the MPC fluid particle velocities $\mathbf{\bm{v}}_{k}$ and positions $\mathbf{\bm{r}}_{k}$ are after a streaming step are given by
@@
which incorporates coupling of monomers dynamics to the background flow. Here, $\mathbf{\bm{F}}_{mon}^{b}$ is the backflow force applied to the each monomer given by Eq. (10) and is solved using the velocity Verlet integration scheme similar to Eq. (12) but using the total force acting on each monomer instead of $F^{b}_{MPC}$ Eq. (10) instead. The SRD collision step now requires the fluid particles and monomers to undergo stochastic rotations relative to the centre of mass velocity of the collision cells which is given by
@@
where $\mathbf{\bm{v}}_{k}(t+\delta t)$ is now determined by Eq. (12) and $\mathbf{\bm{V}}_{i}(t+\delta t)$ is given by the solution of Eq. (13) obtained by the velocity Verlet algorithm. This implies that a Stokeslet flow field by the active force appears for every monomer. The discussed algorithm for external actuated active polymers with tangential active force is implemented in Hoomd-blue Glaser et al. (2015); Anderson et al. (2008) with a home-made modification to include the active and the backflow forces for the coupling of the MPC fluid to the monomer dynamics.
@@ II.4 Dry active polymer simulations
To compare the effects of HI on the polymers, we also carried out Langevin dynamics simulations of dry tangentially-driven polymers described the following equation of motion Isele-Holder et al. (2015) for each monomer $i$:
@@ II.5 Simulation parameters
We choose the monomer diameter $\sigma$ as the unit of length, $k_{B}T$ as the unit of energy and $\tau=\sqrt{m\sigma^{2}/k_{B}T}$ as the unit of time. The wet polymers are simulated in a three-dimensional box with sides of length $L_{box}=60$ and periodic boundary conditions. The simulation box is split into collision cells of size $L_{cell}=1$. The mass of the MPC fluid particles is set to $m=1$ and their collision time step is set to $\delta t=0.01$. The stochastic rotation angle is chosen $\alpha=130^{\circ}$. The number density of the MPC particles is $\langle N_{c}\rangle=10$ leading to a total of $2.16\times 10^{6}$ MPC particles in the simulation box. The viscosity of a solvent made of MPC particles is given by $\eta=82.1\sqrt{\frac{mk_{B}T}{L_{cell}^{4}}}$ Winkler and Huang (2009). Therefore, the chosen parameters create a viscosity of $\eta=82.1$.
We investigate conformation and dynamics of polymers with $N_{m}=50$ monomer of two different values of bending rigidity $\kappa=0$ and $10$ leading to fully flexible and semi-flexible polymers. The excluded volume interaction strength is set to $\epsilon=k_{B}T=1$. We choose the monomer mass to be $M=\langle N_{c}\rangle=10$. The magnitude of the dimensionless active force varies in the range $250.1<f^{a}<$ 25. We set bond rest length $\ell$ to unity and the spring constant of the bonds is chosen as $k_{s}=(10+2f^{a})\times 10^{3}$ to increase linearly with the activity level to prevent bond overstretching and to ensure $\ell=1$. The monomer positions and velocities are updated with time steps of $\Delta t=0.001=\delta t/10$.
We simulate 150 individual wet polymers. For each polymer, we first generate a three dimensional random walk of length $N_{m}$ and step size $\ell$. We use this sequence as the initial positions of the monomers. Once the polymer is created, the system is run for $10^{7}$-$10^{8}$ time steps to reach a steady state depending on the value of $f^{a}$. Then we run a production simulation of $2\times 10^{7}$ $2\times 10^{8}$ time steps, depending on the activity level. Simulations of dry active polymers, conducted for comparison with their wet counterparts, involve 120 individual dry active polymers described by Eq. (15) in a box with size $L_{box}=200$, while all other parameters being identical to those of wet active polymer model. The friction coefficient of monomers in dry active polymer simulation is chosen to be $\gamma=3\pi\eta\sigma\approx 773$ to be compatible with those of wet active polymers. This leads to a ratio of $M/\gamma\approx 0.013$ which ensures that the dynamics of monomers are overdamped and inertial effects  Fazelzadeh et al. (2023) remain negligible. To reach the steady state, we use $10^{7}$ to $10^{9}$ time steps, depending on the activity level. For the production run, twice the number of time steps used to reach the steady state are employed. The time step for dry simulations was set to $\Delta t^{\prime}=0.002$.

### v2

@@ I Introduction
Active particles, driven by self-generated mechanisms, have emerged as a fascinating research area as they exhibit novel patterns of non-equilibrium self-organization with no counterparts in equilibrium systems. Active systems, encompassing a diverse range of entities such as bird flocks, fish schools, swimming bacteria, and artificial microswimmers, offer valuable insights into the governing principles of non-equilibrium self-organization in active matter Vicsek and Zafeiris (2012); Bechinger et al. (2016). Recent research interests have extended beyond point-like or rigid active particles (e.g., active Brownian particles) to include particles with complex shapes or internal degrees of freedom. Prominent examples include active polymers and filaments Winkler et al. (2017); Winkler and Gompper (2020); Zöttl and Stark (2016); Elgeti et al. (2015), which have recently attracted a lot of attention.
@@
Active filaments and polymers can be found on a wide range of scales from ATP-powered chromatin in cell nucleus and motor-driven actin and microtubules to microscopic worms and snakes. Inspired by experiments of active biological filaments a diverse range of theoretical and computational models for active polymers have been proposed in recent years Winkler et al. (2017); Winkler and Gompper (2020) including scalar Smrek and Kremer (2018); Chubak et al. (2020) and force-driven active polymer models where the active force can be stochastic Eisenstecken et al. (2016); Ghosh and Gov (2014), tangential Isele-Holder et al. (2015); Bianco et al. (2018) or perpendicular Prathyusha et al. (2022) to the backbone. Among these, the tangentially-driven active polymer model has garnered increased interest due to its relevance to filament-like systems, such as motor-driven actin and microtubule filaments, elongated bacteria like Proteus mirabilis Lough et al. (2023), and worm species such as T. tubifex  Deblais et al. (2020) that propel themselves along their backbone.
@@
Microscopic active filament-like systems, such as motor-driven biopolymers and bacteria swim in aquatic environments. Consequently, their shape and dynamics are influenced not only by their intrinsic activity mechanisms but also by fluid-mediated interactions. It is well known that the dynamics of passive polymer solutions are significantly affected by hydrodynamic interactions Zimm (1956); Doi et al. (1988); Nikoubashman et al. (2016); Slater et al. (2009). Therefore, it is important to elucidate the interplay between internal degrees of freedom, activity, and hydrodynamic interactions on the conformation and dynamics of active polymers. Integrating shape, activity, and hydrodynamic interactions in studies of active polymer models not only provides a deeper understanding of biophysical processes involving wet active polymers but also inspires the design of artificial filamentous microswimmers with tailored properties.
@@
The majority of studies on active polymers have ignored hydrodynamic interactions and have focused on the interplay between internal dynamics and activity mechanisms. There are a few studies that investigate the role of hydrodynamic interactions on the structural and dynamical properties of active filaments Zöttl and Yeomans (2019); Jayaraman et al. (2012); Laskar and Adhikari (2015); De Canio et al. (2017); Saintillan et al. (2018); Martín-Gómez et al. (2019, 2019); Mahajan and Saintillan (2022); Stein et al. (2021); Lough et al. (2023). These include studies of active Brownian polymers in solution Martín-Gómez et al. (2019, 2019) and active polymer models where monomers are subject to permanent or stochastic active force dipoles along their backbone Jayaraman et al. (2012); Mahajan and Saintillan (2022). These investigations reveal the significance of hydrodynamic interactions on polymer dynamics and even on conformational properties.
@@
In this work, we analyze the conformational and dynamical properties of tangentially-driven active polymers Isele-Holder et al. (2015) embedded in a fluid where the monomers are externally actuated, focusing on unconfined dilute active polymer solutions. The dry version of this model has been subject of intensive studies Isele-Holder et al. (2015); Fazelzadeh et al. (2023); Bianco et al. (2018); Mokhtari and Zippelius (2019); Karan et al. (2024); Khalilian et al. (2024); Duman et al. (2018); Prathyusha et al. (2018); Vatin et al. (2024). This model is relevant for biological filaments actuated by molecular motors and chains of Janus active colloids bonded by electrostatic or magnetic dipolar interactions Yan et al. (2016); Nariaki and C. Patrick (2020); Chao et al. (2024) resulting in an active force tangential to the backbone. Comparing the conformation and dynamics of dry and wet tangentially-driven polymers we address two important questions. (i) Do the hydrodynamic interactions lead to the emergence of new modes of conformation and transport of active polymers? (ii) How do the longtime dynamics of the wet active polymers differ from that of the dry ones?
@@
To account for hydrodynamic interactions, we employ the explicit fluid simulation method of multi particle collision dynamics (MPCD). The MPCD is a mesoscale particle-based simulation method for simulation of fluids, which naturally incorporates thermal fluctuations. The MPCD is able to capture the hydrodynamic correlations correctly Malevanets and Kapral (1999); Howard et al. (2019); Tao et al. (2008); Llahí et al. (2022); Huang et al. (2012) and is widely used for simulating dynamics of colloidal and polymeric solutions. The MPCD method allows to faithfully reproduce the fluid-mediated interactions with significantly less computational costs. In the context of active systems, it has been utilized to study the dynamics of microswimmers such as generic squirmer model Babu and Stark (2012); Hu et al. (2015); Downton and Stark (2009); Götze and Gompper (2010); Zöttl and Stark (2018) and wet active Brownian polymer models Martín-Gómez et al. (2019, 2019). In this article, we present an implementation of wet tangentially-driven polymers in MPCD. The polymer model consists of linear bead-spring chains where each monomer experiences an external active force tangential to the local tangent of the backbone Isele-Holder et al. (2015).
@@
The remainder of this article is organized as follows. In section II, we first introduce our simulation setup including the implementation of tangentially actuated polymer model in the MPCD, providing definition and values of the relevant dimensionless parameters. In Sec. III, We investigate the effects of hydrodynamic interactions (HI) on conformational properties of active polymers. In Sec. IV, we discuss the shrinkage mechanism induced by fluid-mediated interactions. Then, we present the results on impact of hydrodynamic interactions on orientational and translational dynamics of active polymers. We end our paper with a summary of our most important findings and concluding remarks in Sec. V.
@@ II.1 Multiparticle Collision Dynamics simulations
where $\mathbf{\mathcal{R}}(\alpha)$ is a rotation matrix and $\mathbf{\bm{v}}_{cm}=\frac{1}{N_{c}}\sum^{N_{c}}_{i=1}\mathbf{\bm{v}}_{i}$ is the centre of mass velocity of a collision cell. The updated velocities $\mathbf{\bm{v}}^{\prime}_{i}(t+\delta t)$ are then used for repeating the streaming step described by Eq. 1 .
@@ II.2 Active polymer model
We model a semiflexible polymers by a bead-spring chain composed of $N_{m}$ active monomers of mass $M$, with positions $\mathbf{\bm{R}}_{i}$, and velocities $\mathbf{\bm{V}}_{i}$ $(i=1,...,N_{m})$, see Fig. 1. The monomers are connected by harmonic springs described by potential $U_{sp}$. The degree of flexibility of polymers is incorporated by restrictions of bond orientations via a bending potential $U_{bend}$. The self-avoidance is ensured via excluded-volume interactions $U_{EV}$. The dynamics of each monomer is described by the following equation of motion
@@
$U_{EV}(R_{ij})=\begin{cases}4\epsilon\Bigg{[}\bigg{(}\frac{\sigma}{R_{ij}}\bigg{)}^{12}-\bigg{(}\frac{\sigma}{R_{ij}}\bigg{)}^{6}\Bigg{]}+\epsilon\text{,}\;R_{ij}\leq 2^{1/6}\sigma\\ 0\text{,}\quad R_{ij}>2^{1/6}\sigma,\end{cases}$ (7)
where $\sigma$ effectively sets the diameter of the monomers, and $R_{ij}=|\mathbf{\bm{R}}_{i}-\mathbf{\bm{R}}_{j}|$ is the distance between the $i$th and $j$th monomers. The active force $\mathbf{\bm{f}}^{a}_{i}$ in Eq. (3) is an external tangential force directed along the backbone of the polymer mimicking the effect of the tangential force exerted by the molecular motors on the polymer backbone. The active force on each bead, except for the end monomers, is modeled as Isele-Holder et al. (2015):
@@ II.3 Coupling of polymer dynamics to fluid
where $\mathbf{\bm{F}}^{a}_{tot}$ is the sum of all the tangential forces applied to the monomers. The sum of forces induced by intramolecular interactions in the polymer cancel out due to Newton’s third law, thus we can focus purely on the total active force $\mathbf{\bm{F}}^{a}_{tot}$ itself. In confined systems, the global fluid flow creates a backflow. We can achieve a similar backflow in systems with periodic boundary conditions by applying an additional backflow force to all monomers and MPC particles so that the total force becomes null, i.e. $\mathbf{\bm{F}}^{a}_{tot}(t)+\mathbf{\bm{F}}^{b}_{tot}(t)=0$. We distribute the backflow forces acting on the monomers $\mathbf{\bm{F}}^{b}_{mon}$ and MPC particles $\mathbf{\bm{F}}^{b}_{MPC}$ according to their masses Martin-Gomez et al. (2020); Llahí et al. (2022) as
@@
Here, $M$ and $m$ are the masses of the monomers and MPC particles and $N_{m}$ and $N$ denote the number of monomers and MPC particles respectively. The magnitude of the backflow force is governed by the ratio between the particle mass and the total mass of the particles in the system. Assuming that the timescale of changes of polymer conformation are larger than the collision time $\delta t$, the MPC fluid particle velocities $\mathbf{\bm{v}}_{k}$ and positions $\mathbf{\bm{r}}_{k}$ after a streaming step are given by
@@
which incorporates coupling of monomers dynamics to the background flow. Here, $\mathbf{\bm{F}}_{mon}^{b}$ is the backflow force applied to each monomer given by Eq. (10). Eq. 13 is solved using the velocity Verlet integration scheme similar to Eq. (12) but using the total force acting on each monomer instead of $F^{b}_{MPC}$. The SRD collision step now requires the fluid particles and monomers to undergo stochastic rotations relative to the centre of mass velocity of the collision cells which is given by
@@
where $\mathbf{\bm{v}}_{k}(t+\delta t)$ is determined by Eq. (12) and $\mathbf{\bm{V}}_{i}(t+\delta t)$ is given by the solution of Eq. (13) obtained by the velocity Verlet algorithm. This implies that a Stokeslet flow field by the active force appears for every monomer. The discussed algorithm for external actuated active polymers with tangential active force is implemented in Hoomd-blue Glaser et al. (2015); Anderson et al. (2008) with a home-made modification to include the active and the backflow forces for the coupling of the MPC fluid to the monomer dynamics.
@@ II.4 Dry active polymer simulations
To compare the effects of HI on the polymers, we also carried out Langevin dynamics simulations of dry tangentially-driven polymers described by the following equation of motion Isele-Holder et al. (2015). For each monomer $i$ we have:
@@ II.5 Simulation parameters
We choose the monomer diameter $\sigma$ as the unit of length, $k_{B}T$ as the unit of energy and $\tau=\sqrt{m\sigma^{2}/k_{B}T}$ as the unit of time. The wet polymers are simulated in a three-dimensional box with sides of length $L_{box}=60$ and periodic boundary conditions. The simulation box is split into collision cells of size $L_{cell}=1$. The mass of the MPC fluid particles is set to $m=1$ and their collision time step is set to $\delta t=0.01$. The stochastic rotation angle is chosen as $\alpha=130^{\circ}$. The number density of the MPC particles is $\langle N_{c}\rangle=10$ leading to a total of $2.16\times 10^{6}$ MPC particles in the simulation box. The viscosity of the solvent made of MPC particles is given by $\eta=82.1\sqrt{\frac{mk_{B}T}{L_{cell}^{4}}}$ Winkler and Huang (2009). Therefore, the chosen parameters create a viscosity of $\eta=82.1$.
[…]
We simulate 150 individual wet polymers. For each polymer, we first generate a three dimensional random walk of length $N_{m}$ and step size $\ell$. We use this sequence as the initial positions of the monomers. Once the polymer is created, the system is simulated for $10^{7}$-$10^{8}$ time steps to reach a steady state depending on the value of $f^{a}$. Then we run a production simulation of $2\times 10^{7}$ $2\times 10^{8}$ time steps, depending on the activity level. Simulations of dry active polymers, conducted for comparison with their wet counterparts, involve 120 individual dry active polymers described by Eq. (15) in a box with size $L_{box}=200$, while all other parameters being identical to those of wet active polymer model. The friction coefficient of monomers in dry active polymer simulation is chosen to be $\gamma=3\pi\eta\sigma\approx 773$. This leads to a ratio of $M/\gamma\approx 0.013$ which ensures that the dynamics of monomers are overdamped and inertial effects Fazelzadeh et al. (2023) remain negligible. To reach the steady state, we use $10^{7}$ to $10^{9}$ time steps, depending on the activity level. For the production run, twice the number of time steps used to reach the steady state are used. The time step for dry simulations is $\Delta t^{\prime}=0.002$.